# ParallelComp: Parallel Long-Context Compressor for Length Extrapolation

### v1

@@ 5.2 Length Extrapolation Settings
Main Results
We present our method in Table 1, showing the performance of several strong baselines on LongBench. We have the following main findings: i): Our method is the only one that surpasses FullKV (i.e., the baseline without any length extrapolation) across different backbones. ii): Section 4 reveals that parallel KV cache compression exacerbates attention bias. However, combining it with the eviction $R_{h}$ method to calibrate the attention distribution, i.e., Ours-calibration-compression, can significantly restore the performance to that of the original KV cache size. iii): Chunk-based length extrapolation methods, such as InfLLM and ChunkLlama, generally perform better than position encoding-based methods such as Dynamic-PI and NTK-Aware. iv): Directly calibrating the attention distribution for NTK-aware length extrapolation methods, namely AttenCalibration-NTK, achieves good performance only on longest datasets, such as NtrvQA, GovReport, and RB-P.
@@
We present the results of perplexity (PPL) calculations on the NarrativeQA test set in Table 3, which reflect the model’s performance in long-context language modeling. i): Chunk-based position extrapolation methods (ChunkLlama, InfLLM, and Ours) achieve significantly lower PPL compared to position encoding-based methods (Dynamic-PI and NTK-Aware). ii): Position encoding-based methods start to collapse in performance for language modeling when the length exceeds 32k. iii): As the number of chunks increases (from 2K to 128K), our method still demonstrates consistent perplexity stability across different lengths.
@@
Llama2-7B-chat-hf(4k) Methods R.PK R.Num R.KV En.MC Math.F Code.Debug Average Max Length 125k 125k 175k 834k 120k 258k 273k Ours 100.00 97.63 20.60 33.62 19.71 25.13 49.45 Ours-calibration 100.00 98.64 30.20 36.24 19.71 31.47 52.71 Sink-eviction-layer-1-8 54.24 42.37 4.60 33.62 18.86 25.13 29.80 Sink-eviction-layer-9-16 89.83 81.36 14.60 34.50 18.00 30.66 44.82 Sink-eviction-layer-17-23 98.81 98.31 20.40 32.75 18.86 29.87 49.83 Sink-eviction-layer-24-31 98.98 97.29 20.60 36.24 17.71 31.23 50.34 Recency-eviction-layer-1-8 98.47 97.63 22.40 31.12 19.71 28.98 49.72 Recency-eviction-layer-9-16 98.98 97.97 25.60 32.75 19.71 31.23 51.04 Recency-eviction-layer-17-23 99.15 97.97 27.60 35.20 20.29 21.47 50.28 Recency-eviction-layer-24-31 98.98 97.97 29.80 37.33 19.71 22.38 51.03 Middle-eviction-layer-1-8 98.64 98.31 24.80 38.46 19.14 25.38 50.79 Middle-eviction-layer-9-16 99.32 98.31 23.20 36.24 19.43 30.21 51.12 Middle-eviction-layer-17-23 98.47 98.47 28.40 35.12 19.71 29.45 51.61 Middle-eviction-layer-24-31 99.15 98.14 30.00 31.01 19.71 27.95 50.99
@@ 5.3 Ablation of Attention Bias
We present in Table 4 the impact of evicting different bias tokens at various layers on different tasks. We have the following observations: i): The $R_{s}$ in the shallow layers (1-8) is crucial for retrieval tasks. Without these tokens, the model’s performance will be significantly impaired. ii): The $R_{r}$ in the deeper layers (layers 17-24 and 25-32) is crucial for the model’s coding abilities. Evicting these tokens will lead to a decline in coding performance. iii): Shallow $R_{m}$ (layers 1-8) damages the model’s understanding ability, and evicting them can improve the model’s performance. iv): $R_{r}$ in the early layers (layers 1-8) is important for the model’s in-context learning ability. For a detailed analysis of this phenomenon, please refer to Appendix B.
Chunk Number Ours Ours-compression prefill(s) / generation(ms/token) max memory used(MB) prefill(s) / generation(ms/token) max memory used(MB) 1 1317.72 / 24.30 19394 1317.72 / 22.16 16994 4 321.40 / 43.69 35518 321.40 / 23.28 24734 8 160.70 / 72.53 47758 160.70 / 31.21 36396 12 111.67 / 102.94 65980 111.67 / 39.61 48458 16 - Out-of-Memory 82.36 / 49.25 59140 20 - Out-of-Memory 65.23 / 56.25 71302 23 - Out-of-Memory 56.07 / 57.19 79742 24 - Out-of-Memory - Out-of-Memory
Table 5: Throughput analysis. We evaluate on Llama2-7B-chat-hf and compare the improvement in chunk throughput with the use of parallel KV cache compression. Time tests were performed on the narrativeqa dataset.
@@ 6 Conclusion
In this paper, we propose ParallelComp, a training-free approach to enhance the long-context extrapolation capability of LLMs. By leveraging parallel attention mechanisms and an efficient compression strategy, our method effectively mitigates attention bias while maintaining high efficiency. Experimental results on LongBench and InfiniteBench show that ParallelComp achieves strong performance, surpassing existing extrapolation methods and handling context lengths up to 128K tokens. Notably, our approach achieves 91.17% of GPT-4’s performance on ultra-long context tasks with an 8B model, demonstrating its effectiveness with lower memory overhead. ParallelComp provides a scalable and efficient solution to training-free extension of context lengths in LLMs, enabling lossless improvement in throughput while reducing latency during the prefill stage.
@@
Appendix A Proof for Parallel Attention Collapse

### v2

@@ 5.2 Length Extrapolation Settings
Main results.
We present our method in Table 1, showing the performance of several strong baselines on LongBench. We have the following main findings: i): Our method is the only one that surpasses FullKV (i.e., the baseline without any length extrapolation) across different backbones. ii): Section 4 reveals that parallel KV cache compression exacerbates attention bias. However, combining it with the eviction $R_{H}$ method to calibrate the attention distribution, i.e., Ours-calibration-compression, can restore the performance to that of the original KV cache size. iii): Chunk-based length extrapolation methods, such as InfLLM and ChunkLlama, generally perform better than position encoding-based methods such as Dynamic-PI and NTK-Aware. iv): Directly calibrating the attention distribution in NTK-aware length extrapolation methods, such as AttenCalibration-NTK, leads to strong performance primarily on the longest datasets, including NtrvQA, GovReport, and RB-P. This suggests that the effect of attention distribution calibration becomes increasingly significant as input length grows.
@@
To further compare the performance of our method in language modeling, we present the results of perplexity (PPL) calculations on the NarrativeQA test set in Table 3, which reflect the model’s performance in long-context language modeling. For fair comparison, we typically calculate the PPL for the query chunk, as it corresponds to the model’s decoding phase. i): Chunk-based position extrapolation methods (ChunkLlama, InfLLM, and Ours) achieve significantly lower PPL compared to position encoding-based methods (Dynamic-PI and NTK-Aware). ii): Position encoding-based methods start to collapse in performance for language modeling when the length exceeds 32k. iii): As the number of chunks increases (from 2K to 128K), our method still demonstrates consistent perplexity stability across different lengths. Surprisingly, ChunkLlama maintains high performance on Llama3-8B-instruct, outperforming other methods.
Method NARR QAS MULT HOPT 2WKI MUS GOV QMS NEWS TREC TRIV SSM PCNNT PREN LCC REP AVG
Llama3.1 27.80 44.25 49.46 47.86 40.54 23.64 32.64 22.90 26.90 38.00 88.44 25.64 2.02 92.00 10.35 18.64 36.94
ParallelComp-Llama3.1 29.45 45.98 50.67 48.36 46.56 23.32 32.60 24.29 27.34 38.50 86.72 25.93 0.05 95.00 14.15 21.42 38.15
Qwen2.5 27.83 41.31 50.41 53.52 44.68 30.00 33.38 24.01 25.40 71.00 86.10 39.91 7.25 100.00 6.86 7.88 40.60
ParallelComp-Qwen2.5 28.42 42.24 50.54 56.26 42.02 28.25 33.43 23.20 25.20 71.50 89.21 41.84 5.00 93.50 20.73 13.34 41.54
Table 4: Performance on LongBench benchmark. Models are evaluated under a 24K KV cache budget.
Method PS NUM KV EN.MC MATH CODE AVG
Llama3.1 5.59 26.25 18.60 32.86 31.52 22.56 26.36
ParallelComp-Llama3.1 100.00 83.56 88.60 66.38 37.14 22.08 59.55
Qwen2.5 59.32 58.31 33.80 61.39 85.71 23.76 53.72
ParallelComp-Qwen2.5 100.00 76.27 63.40 66.86 92.57 24.75 70.64
Table 5: Performance on InfiniteBench with different models. Models are evaluated under a 24K KV cache budget.
5.3 Evaluation in Long-context Models
To demonstrate the effectiveness of our method for long-context models, we evaluate it on two models trained with contexts up to 128K, which therefore do not require any extrapolation capabilities: LLaMA3.1 (Grattafiori et al., 2024) and Qwen2.5 (Yang et al., 2025a), used as base models. All evaluations are conducted under a consistent 24K KV cache size. While standard baselines directly utilize the full 24K position encodings, our method applies extrapolation techniques to reuse a 6K position encoding, thereby supporting longer input lengths without modifying the model architecture. To ensure fairness, we standardize the use of special prompt tokens across models (e.g., <|begin_of_text|> for LLaMA3.1), which we observe to have a significant impact on performance. All reported results correspond to configurations that include these prompt tokens.
As shown in Table 4 and Table 5, our ParallelComp method consistently improves performance across a broad range of tasks. On LongBench, we observe modest but consistent gains compared to baselines. On InfiniteBench, which emphasizes ultral-long context understanding, our method demonstrates significant improvements, particularly for tasks such as PS, NUM, and KV. These results indicate that extrapolated position encodings can be effectively reused in parts of the model’s position encodings to extend context length, achieving comparable or even improved performance relative to using full position encodings in long-context models.
@@ 5.3 Evaluation in Long-context Models
Llama2-7B-chat-hf(4k) Methods R.PK R.Num R.KV En.MC Math.F Code.Debug Average Max Length 125k 125k 175k 834k 120k 258k 273k Ours 100.00 97.63 20.60 33.62 19.71 25.13 49.45 Ours-calibration 100.00 98.64 22.80 36.24 19.71 30.20 51.27 Sink-eviction-layer-1-8 99.32 42.71 2.20 37.12 17.71 22.84 36.98 Sink-eviction-layer-9-16 100.00 91.19 11.00 37.12 14.86 24.37 46.42 Sink-eviction-layer-17-23 100.00 97.80 20.80 33.19 19.14 30.96 50.32 Sink-eviction-layer-24-31 100.00 97.63 20.20 31.88 18.00 29.19 49.48 Recency-eviction-layer-1-8 100.00 96.44 2.60 33.19 16.00 19.54 44.63 Recency-eviction-layer-9-16 100.00 97.80 15.80 37.99 10.86 23.10 47.59 Recency-eviction-layer-17-23 100.00 97.97 20.40 23.58 16.00 32.74 48.45 Recency-eviction-layer-24-31 100.00 97.63 20.60 35.81 18.57 25.89 49.75 Middle-eviction-layer-1-8 100.00 97.29 20.60 34.93 18.29 22.84 48.99 Middle-eviction-layer-9-16 100.00 97.63 20.60 33.62 16.00 26.40 49.04 Middle-eviction-layer-17-23 98.81 97.46 20.00 34.06 19.14 30.20 49.95 Middle-eviction-layer-24-31 100.00 97.46 19.80 30.13 19.43 28.93 49.29
@@ 5.4 Ablation of Attention Bias
We present in Table 6 the impact of evicting different bias tokens at various layers on different tasks. We have the following observations: i): The $R_{s}$ in the shallow layers (1-8) is crucial for retrieval tasks. Without these tokens, the model’s performance will be significantly impaired. ii): The $R_{r}$ in the deeper layers (layers 9-16) plays a crucial role in the model’s reasoning abilities. Evicting these tokens results in a decline in performance on coding and math tasks. iii): Shallow $R_{m}$ (layers 1-8) damages the model’s understanding ability, and evicting them can improve the model’s performance. Deep $R_{m}$ (layers 24-31) contributes to the model’s ability in reading comprehension tasks (En.MC), and evicting them harms the model’s performance. iv): $R_{r}$ in the early layers (layers 1-8) is important for the model’s in-context learning ability. For a detailed analysis of this phenomenon, please refer to Appendix B.
Chunk Number Ours Ours-Compression Prefill (s) / Generation (ms/token) Max Memory Used (MB) Prefill (s) / Generation (ms/token) Max Memory Used (MB) 1 1317.72 / 24.30 19394 1317.72 / 22.16 16994 4 321.40 / 43.69 35518 321.40 / 23.28 24734 8 160.70 / 72.53 47758 160.70 / 31.21 36396 12 111.67 / 102.94 65980 111.67 / 39.61 48458 16 N/A Out-of-Memory 82.36 / 49.25 59140 20 N/A Out-of-Memory 65.23 / 56.25 71302 23 N/A Out-of-Memory 56.07 / 57.19 79742 24 N/A Out-of-Memory N/A Out-of-Memory
Table 7: Throughput analysis. We evaluate on Llama2-7B-chat-hf and compare the improvement in chunk throughput with the use of parallel KV cache compression. Time tests were performed on the NarrativeQA dataset. Experiments are conducted on an AMD Instinct MI210 64GB GPU.
@@ 6 Conclusion
In this paper, we propose ParallelComp, a training-free and parallel long-context compression framework that significantly enhances the extrapolation capability of large language models (LLMs) for ultra-long contexts. ParallelComp overcomes the critical memory bottlenecks in length extrapolation and systematically analyzes the attention bias that arises in such settings. Specifically, our method allows 8B LLMs to extend inference length from 4K to 128K tokens on a single A100 80GB GPU without retraining or significant degradation in performance. By leveraging chunk-based parallel attention, dynamic KV cache eviction, and an attention calibration strategy, our approach alleviates both excessive memory usage and the attention sink phenomenon. Extensive theoretical and empirical results demonstrate that ParallelComp effectively mitigates attention bias and enables robust, end-to-end inference. Notably, our method achieves $91.17\%$ of GPT-4’s performance on ultra-long context tasks using an 8B model, outperforming various state-of-the-art closed-source models. These findings pave the way for more scalable and efficient long-context inference.
@@
Acknowledgements
This work is supported in part by the Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council (RGC), Hong Kong SAR, and in part by the AVNET-HKU Emerging Microelectronics & Ubiquitous Systems (EMUS) Lab.
Future Work
While ParallelComp enables efficient length extrapolation up to 128K tokens, future work can explore its application to long-chain reasoning and ultra-long in-context learning (Xiong et al., 2023). This includes designing methods for reasoning over 100K+ tokens, developing test-time scaling. Further exploration into in-context learning will also be critical for reliable deployment in real-world long-context reasoning.
Appendix A Proofs for Parallel Attention Bias
A.1 Exponential Decay of Local Attention
We formalize the exponential decay of attention values as a function of relative distance when using relative position encodings (Su et al., 2024). This behavior is modulated by a sparsity control parameter $R$, which regulates the rate at which attention mass diminishes over increasing distance.
Theorem A.1 (Exponential Decay of Attention).
Let $A^{c}_{\mathfrak{l}}[i,j]$ denote the local attention weight from token $i$ to token $j$ within chunk $c$, computed using scaled dot-product attention with relative position encoding. Suppose the dot product between query $q_{i}$ and key $k_{j}$ depends linearly on their relative distance, i.e.,
$q_{i}\cdot k_{j}\approx-\alpha\cdot\text{d}(i,j),$
where $\text{d}(i,j)$ is the positional distance and $\alpha>0$ is a constant. Then the attention value exhibits exponential decay:
$A^{c}_{\mathfrak{l}}[i,j]\sim\exp\left(-O(R)\cdot\text{d}(i,j)\right),$
where $R$ is a sparsity parameter controlling the effective attention range.
Proof.
The local attention matrix is given by
$A^{c}_{\mathfrak{l}}[i,j]=\frac{\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}\right)}{\sum_{j^{\prime}=1}^{w}\exp\left(\frac{q_{i}\cdot k_{j^{\prime}}}{\sqrt{d}}\right)}.$
Assuming $q_{i}\cdot k_{j}\approx-\alpha\cdot\text{d}(i,j)$, we obtain:
$\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}\right)\approx\exp\left(-\frac{\alpha}{\sqrt{d}}\cdot\text{d}(i,j)\right).$
Letting $O(R):=\alpha/\sqrt{d}$ leads to:
$A^{c}_{\mathfrak{l}}[i,j]\propto\exp\left(-O(R)\cdot\text{d}(i,j)\right),$
establishing the exponential decay behavior. ∎
Discussion.
This exponential decay arises from the softmax mechanism and relative position encoding: as the relative distance increases, the inner product diminishes, leading to exponentially suppressed attention values. The parameter $R$ determines how rapidly this suppression occurs. A higher $R$ yields slower decay (broader attention), while a lower $R$ enforces sharper locality—crucial for analyzing sparsity and receptive fields in attention models.
Theorem A.2 (Tail Bound on Effective Attention Mass).
Suppose $A^{c}_{\mathfrak{l}}[i,j]\sim\exp(-O(R)\cdot\text{d}(i,j))$ with $R=O(\sqrt{\log w})$. Let $\epsilon>0$ be a threshold and $k$ be the number of ineffective entries in row $i$ of the local attention matrix, i.e., those satisfying $A^{c}_{\mathfrak{l}}[i,j]\leq\epsilon$. Then for any $10<\delta<$ 1,
$\epsilon\geq\exp\left(-O(R)\cdot\sqrt{\log\left(\frac{w(w-k)}{\delta}\right)}\right).$
Proof.
Let us consider the attention weight between query $q_{i}$ and key $k_{j}$ (within chunk $c$) given by
$A^{c}_{\mathfrak{l}}[i,j]=\frac{\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}\right)}{\sum_{j^{\prime}=1}^{w}\exp\left(\frac{q_{i}\cdot k_{j^{\prime}}}{\sqrt{d}}\right)}.$
Assume the dot product $q_{i}\cdot k_{j}$ (after position encoding) satisfies
$q_{i}\cdot k_{j}=-\alpha\cdot d(i,j)+Z_{i,j},$
where $d(i,j)$ is the relative distance, according to Theorem A.1, $\alpha>0$ is a scaling constant, and $Z_{i,j}$ is a mean-zero, sub-Gaussian noise term arising from randomness in $q_{i}$ and $k_{j}$.
Sub-Gaussian concentration.
Suppose $Z_{i,j}$ is sub-Gaussian with parameter $\sigma^{2}$ (variance proxy). Since $q_{i},k_{j}\in\mathbb{R}^{d}$ are assumed to be independent and isotropic (or sufficiently well-behaved), the inner product $Z_{i,j}$ scales as $\sim\mathcal{N}(0,d)$, so after normalization by $\sqrt{d}$, we have:
$\frac{Z_{i,j}}{\sqrt{d}}\sim\mathcal{N}(0,1)$
and more generally, sub-Gaussian(1) for all entries.
Tail probability for an entry exceeding $\epsilon$.
We are interested in the probability a particular attention entry $A^{c}_{\mathfrak{l}}[i,j]$ exceeds a fixed threshold $\epsilon>0$:
$\mathbb{P}\left(A^{c}_{\mathfrak{l}}[i,j]>\epsilon\right).$
The numerator is $\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}\right)$, which, via the deterministic and random parts, is $\exp\left(-\frac{\alpha}{\sqrt{d}}\text{d}(i,j)+\frac{Z_{i,j}}{\sqrt{d}}\right)$. The denominator $Z_{i}$ is the sum over all $w$ such terms.
By the softmax property, after normalization, it holds that for entries with large $d(i,j)$ (i.e., far from the diagonal), the probability concentrates at low values, and for an entry to exceed $\epsilon$ the exponent must be large:
$\exp\left(-\frac{\alpha}{\sqrt{d}}d(i,j)+\frac{Z_{i,j}}{\sqrt{d}}\right)>\epsilon Z_{i}$
or
$-\frac{\alpha}{\sqrt{d}}d(i,j)+\frac{Z_{i,j}}{\sqrt{d}}>\log(\epsilon Z_{i}).$
Union bound over all entries.
We need to ensure with high probability (at least $1-\delta$) that at most $k$ entries per row (out of $w$) exceed $\epsilon$. That is, we want (by union bound)
$w(w-k)\cdot\mathbb{P}\left(A^{c}_{\mathfrak{l}}[i,j]>\epsilon\right)\leq\delta.$
Then,
$\mathbb{P}\left(A^{c}_{\mathfrak{l}}[i,j]>\epsilon\right)\leq\frac{\delta}{w(w-k)}$
which, after inverting the softmax and using sub-Gaussian tail estimates for $Z_{i,j}$, gives
$\mathbb{P}\left(\frac{Z_{i,j}}{\sqrt{d}}>\log(\epsilon Z_{i})+\frac{\alpha}{\sqrt{d}}\text{d}(i,j)\right)\leq\frac{\delta}{w(w-k)}.$
The sub-Gaussian tail bound reads for $t>0$:
$\mathbb{P}(X>t)\leq\exp(-t^{2}/2).$
Setting $t^{2}/2=\log\left(\frac{w(w-k)}{\delta}\right)$ leads to
$t=\sqrt{2\log\left(\frac{w(w-k)}{\delta}\right)}$
and relating $t$ back to normalization and decay constants, finally yields
$\epsilon\gtrsim\exp\left(-O(R)\cdot\sqrt{\log\left(\frac{w(w-k)}{\delta}\right)}\right)$
where $R=\alpha/\sqrt{d}$ (showing dependence on $d$).
The role of $d$ in the bound.
The normalization by $\sqrt{d}$ plays a key role: as $d$ increases, the effective noise in the dot product becomes less significant relative to the deterministic decay $-\alpha d(i,j)$, pulling most attention values closer to zero except for a small set (nearby positions). This effect concentrates the attention and tightens the tail, making large attention far from the diagonal even less probable.
Note that, in total union bound, we consider $w^{2}$ possible entries and sometimes further divide by $d$, reflecting the increased concentration with high dimension or the independence among $d$ channels (as found in certain information-theoretic arguments for high-dimensional softmax). This dependence enters the denominator for the upper bound on $k$:
$k\leq w-\exp\left(O\left(\frac{\log^{2}(\epsilon w)}{R^{2}}\right)\right)\frac{\delta}{wd}.$
Summary.  The dependence on $d$ in the denominator comes from sub-Gaussian concentration for scalar products in dimension $d$. As $d$ grows, the likelihood of far-off entries (large distance) carrying significant attention decays rapidly due to stronger concentration of measure, resulting in increased sparsity of the attention matrix for fixed threshold $\epsilon$. Therefore, for sufficiently large $w$ and $d$, with fixed $\epsilon$,
$k\ll w,$
implying
where $k=o(w)$.
Remark.
The appearance of the term $d$ in the denominator reflects how higher model dimensionality increases the concentration of dot products, making rare large excursions (that could create significant attention values far from the main focus) even less probable. Thus, the sparsity bound becomes stricter as $d$ increases. This probabilistic tail behavior is essential for understanding how the structure and dimension of the attention model affect the sparsity and locality of its learned representations.
A.2 Sparsity in Local Attention
Theorem A.3 (Sparsity of Local Attention).
Let $X\in\mathbb{R}^{N\times d}$ be partitioned into $C$ chunks of width $w$. For each chunk $c$, let local attention matrix $A^{c}_{\mathfrak{l}}\in\mathbb{R}^{w\times w}$ be computed via
$A^{c}_{\mathfrak{l}}=\mathrm{Softmax}\left(\frac{Q^{c}{K^{c}}^{\top}}{\sqrt{d}}\right).$
Let $\mathcal{S}_{\epsilon}^{(c)}(A^{c}_{\mathfrak{l}}[i,:])=\{j\in[w]~{}|~{}A^{c}_{\mathfrak{l}}[i,j]>\epsilon\}$ for fixed $\epsilon>0$, and set $R=O(\sqrt{\log w})$.
Then for any $10<\delta<$ 1, as $w\to\infty$, with high probability $1-\delta$,
$\left|\mathcal{S}_{\epsilon}^{(c)}(A^{c}_{\mathfrak{l}}[i,:])\right|=w-k,$
where $k=o(w)$ depends on $\epsilon$ and $R$.
Proof.
We begin with Theorem A.2, which provides a tail bound for the effective attention mass:
$\epsilon\geq\exp\left(-O(R)\cdot\sqrt{\log\left(\frac{w(w-k)}{\delta}\right)}\right).$
Taking logarithms on both sides gives
$\log\epsilon\geq-O(R)\cdot\sqrt{\log\left(\frac{w(w-k)}{\delta}\right)}.$
Rearranging, we have
$\sqrt{\log\left(\frac{w(w-k)}{\delta}\right)}\leq\frac{|\log\epsilon|}{O(R)},$
so
$\log\left(\frac{w(w-k)}{\delta}\right)\leq\frac{\log^{2}\epsilon}{O(R^{2})}.$
Exponentiating both sides, we have
$\frac{w(w-k)}{\delta}\leq\exp\left(O\left(\frac{\log^{2}\epsilon}{R^{2}}\right)\right),$
which implies
$w(w-k)\leq\delta\exp\left(O\left(\frac{\log^{2}\epsilon}{R^{2}}\right)\right).$
Solving for $k$ gives
$k\geq w-\frac{1}{w}\delta\exp\left(O\left(\frac{\log^{2}\epsilon}{R^{2}}\right)\right).$
Now, if we refine the union bound to account for all entries in $w\times w$ attention matrix for $d$-dimensional vectors, the probability of an entry exceeding $\epsilon$ is further divided by $d$ due to the independence in $d$ dimensions (assuming the dot products across dimensions act as independent sub-Gaussian random variables). That is, the standard deviation of the normalized dot product is reduced by a factor of $\sqrt{d}$, affecting the spread of attention weights. Thus, the probability becomes
$\mathbb{P}(A^{c}_{\mathfrak{l}}[i,j]>\epsilon)\leq\frac{\delta}{w(w-k)d},$
leading to the sparsity bound
$k\leq w-\exp\left(O\left(\frac{\log^{2}(\epsilon w)}{R^{2}}\right)\right)\frac{\delta}{wd}.$
The presence of $d$ in the denominator reflects how higher model dimensionality increases concentration, making the tail events (large attention weights far from the focus) less probable.
Therefore, for sufficiently large $w$ and fixed $\epsilon$, we indeed have $k\ll w$, and thus
with $k=o(w)$. ∎
Remark.
The inclusion of the $d$ term in the denominator arises from the probabilistic tail bound under the assumption that the dot products are sub-Gaussian in $d$ dimensions. As $d$ increases, the attention weights become more concentrated (due to the central limit effect over the $d$ independent dimensions), reducing the portion of entries that exceed a fixed threshold $\epsilon$. This effect is captured in the sparsity upper bound through the explicit dependence on $d$.
A.3 Exponential Decay in Global Attention
We extend the above result to global attention, which spans the entire sequence.
Theorem A.4 (Exponential Decay of Global Attention).
Let $A_{\mathfrak{g}}[i,j]$ be the global attention from query $i$ to key $j$, where
$q_{i}\cdot k_{j}\approx-\alpha\cdot\text{d}(i,j).$
Then,
$A_{\mathfrak{g}}[i,j]\sim\exp\left(-O(R)\cdot\text{d}(i,j)\right),$
with $R=\alpha/\sqrt{d}$.
Proof.
Analogously to the local case, the attention value is
$A_{\mathfrak{g}}[i,j]\approx\frac{\exp\left(-\frac{\alpha}{\sqrt{d}}\text{d}(i,j)\right)}{Z_{i}}\propto\exp\left(-O(R)\text{d}(i,j)\right),$
where $Z_{i}$ is the row normalization. ∎
A.4 Tail Bound for Global Attention
Theorem A.5 (Tail Bound for Global Attention Mass).
Suppose $A_{\mathfrak{g}}[i,j]\sim\exp(-O(R)\cdot\text{d}(i,j))$. Let $k_{\mathfrak{g}}$ denote the number of global attention entries below $\epsilon$ in each row. Then
$\epsilon\geq\exp\left(-O(R)\cdot\sqrt{\log\left(\frac{(Cw+w_{q})(w_{q}-k_{\mathfrak{g}})}{\delta}\right)}\right).$
Proof.
Applying the union bound and exponential decay form as in the local case,
$\mathbb{P}(A_{\mathfrak{g}}[i,j]>\epsilon)\leq\frac{\delta}{(Cw+w_{q})(w_{q}-k_{\mathfrak{g}})},$
which yields the stated bound for $\epsilon$. ∎
A.5 Sparsity of Global Attention